OpenStudy (anonymous):

Jeremie ran around the track at the YMCA for 2 hours. When he was done he figured that he had traveled 20 kilometers. What was his speed? What was his velocity?

OpenStudy (anonymous):

Speed= distance/time = 20/2 = 10 kmph Velocity= displacement/time =0/2 = 0 kmph #in case you want the answer in ms^-1, multiply the answer by 5/18